will you please help me with this equation? im having a hard time figuring it out.
Solve the quadratic equation
1/1+x-1/3-x=6/35

Answer by stanbon(29365)

1/(x+1) + 1/(x-3) = 6/35
Multiply thru by 35(x+1)(x-3) to get:
35(x-3) + 35(x+1) = 6(x^2-2x-3)
70x-70 = 6(x^2-2x-3)
35x-35 = 3(x^2-2x-3)
3x^2-41x+26=0
Use the quadratic formula to get:
x=[41+-sqrt(41^2-4*3*26]/6
x=[41+-37]/6
x=(41+37)/6 or x=(41-37)/6
x=78/6 or x=4/6
x=38/3 or x=2/3
--------
Cheers,
Stan H.